OpenStudy (ray_sihota):

Need help Yolanda is due to make a payment of a 1000 now. instead she negotiated to make two equal payments one year and two years from now . Determine the size of equal payments if money is worth 8% compounded quarterly. Use now as the focal date

OpenStudy (ray_sihota):



OpenStudy (anonymous):

Ok I have an answer .... But I used the TMV Solve is that alright?

OpenStudy (anonymous):

or do you need to use an equation?

OpenStudy (ray_sihota):

tvm

OpenStudy (dumbcow):

its a present value problem \[PV = P(v +v^2) = 1000\] \[P = \frac{1000}{v+v^2}\] where \[v = \frac{1}{1+i} = \frac{1}{1.02^4}\]

OpenStudy (anonymous):

@dumbcow hmmm I was thinking of approaching it diff But either way do you guys know how to work it out that interest is compounded quarterly and payment annualy ... still tryna figure out this TVM solver

OpenStudy (dumbcow):

@wahahaha , you have to adjust the interest to annual annual effective interest is (1.02)^4 - 1

OpenStudy (anonymous):

ohhh dang it ... I shld have remembered abt AER

OpenStudy (ray_sihota):

this is what I got so far. Because I going past the focal date I did FV = x(1+ 0.08/4) expoent -4x1 and 2nd payment is x(1+0.08/4) exponent -4x2= x(1.02 )expnant -8 and -4 x1 x(1.02) exponent -4 i am lost after that

OpenStudy (anonymous):

Hey its hard to follow what u r saying ... Hey maybe you wanna use the drawing option

OpenStudy (ray_sihota):

gimme a minute

OpenStudy (anonymous):

\(x(1+\frac{.08}{4})^{-4*1}+x(1+\frac{.08}{4})^{-4*2}=x(1.02)^{-8}+x(1.02)^{-4}\) Is that what u r tryna say?

OpenStudy (ray_sihota):

yes

OpenStudy (ray_sihota):

that is how my professor showed me but i lost him , after that step

OpenStudy (dumbcow):

thats what i showed, sorry my notation is different here is a reference for my notation: https://en.wikipedia.org/wiki/Present_value#Present_value_of_a_lump_sum

OpenStudy (ray_sihota):

if my payments go past the focal date how can they be present value?

OpenStudy (anonymous):

Ok i just use TVM solver and got equal payments of $562.62 ...

OpenStudy (dumbcow):

@ray_sihota , factor out x, then divide \[x = \frac{1000}{1.02^{-4} + 1.02^{-8}}\]

OpenStudy (dumbcow):

no the payments are made in the future, but the original 1000 is in the present, thus the present value of all future payments

OpenStudy (anonymous):

Ya we have a PV of 1000 bucks which has an interest of 8% compounded annually so all u r trying to do is find the PMT I entered it into TVM as follows N=2 I/Y=Effective Annual Rate = 8.24 PV=1000 Then computer Pmt and your answer will be 562.62

OpenStudy (dumbcow):

yep ^

OpenStudy (anonymous):

@dumbcow Do you mind explaining the ray's method I only study using TVM never learn abt equations unfortunately

OpenStudy (ray_sihota):

got it thank you so much , this explanation makes so much sense.

OpenStudy (dumbcow):

:)